# No Verifiable Reward for Prosody: Toward Preference-Guided Prosody Learning in TTS

### v1

@@ 1 Introduction
To close this reward gap, we adopt iterative Direct Preference Optimization (DPO) with small human-in-the-loop batches (200 preference pairs per round). Across rounds, DPO supplies a directly verifiable signal for prosodic naturalness while regularizing to the current model, yielding both improved human preference and competitive CER (Fig. 1). Our contributions are as follows:
@@
  • We release KoCC-TTS, an evaluation set of frequent Korean call-center phrases for TTS prosody assessment.
@@ 2 Related Work
GRPO for TTS and reward design Recent TTS studies that adopt group-relative policy optimization (GRPO) primarily target intelligibility and identity preservation by rewarding ASR-derived errors and speaker similarity, sometimes adding non-intrusive quality predictors. For instance, F5R-TTS couples WER with speaker-similarity (SIM) under GRPO [9], DMOSpeech2 optimizes a duration policy with SIM+WER via GRPO [6], and the TTS-1 technical report describes a composite GRPO reward that blends WER/SIM/DNSMOS for RL alignment [7]. While these report lower error rates and stronger speaker faithfulness, they largely omit explicit prosody-sensitive rewards (e.g., pitch movement, phrasing, boundary control). In practice, we also observed that CER/NLL-oriented GRPO can collapse prosodic variation into near-monotone renderings, consistent with reports of punctuation- and phrasing-related failures in state-of-the-art systems [5].
Preference-based objectives for prosody Direct Preference Optimization (DPO) offers a complementary route by optimizing pairwise human (or proxy) preferences without training a separate reward model [8]. In TTS, Emo-DPO applies DPO to better capture subtle emotional/prosodic nuances with an LLM-TTS backbone, improving both prosody similarity and perceived naturalness [10]. Beyond single-shot DPO, iterative preference optimization for speech generation has also been explored: SpeechAlign constructs codec-token preference pairs and refines a speech LM in multiple rounds, demonstrating iterative self-improvement [11]. Concurrently, differentiable or multi-dimensional preference objectives have been proposed to move past coarse ASR metrics—e.g., DiffRO directly optimizes differentiable rewards over codec tokens [12], and MPO considers multi-criteria screening of preference pairs in speech synthesis [13]. These studies collectively suggest that preference-based post-training is a promising way to recover communicative prosody without sacrificing robustness to transcription errors.
@@ 4.2 Setup
We evaluate 12 systems on the KoCC-TTS dataset, including 3 production-grade external services, 3 open-source models, and 6 internal variants. We intentionally exclude open-source TTS baselines from the main comparison. A preliminary screening indicated that most off-the-shelf OSS voices exhibited inadequate prosodic fluency in Korean, and their inclusion would likely result in floor effects rather than provide a meaningful basis for comparison. For external systems, we adopt vendors’ strongest Korean voices and default synthesis settings unless otherwise noted: ElevenLabs Multilingual v2 (“Anna”) [^3] https://elevenlabs.io/blog/eleven-multilingual-v2, Supertone (“sona_speech_1”) [^4] https://docs.supertoneapi.com/ko/user-guide/quickstart, and GPT-4o-mini-tts (“sage”). To ensure fairness, all systems synthesize from the same prompts with identical text normalization rules; speaking rate and punctuation handling are held fixed, and outputs are evaluated at vendors’ native sampling configurations. We report (i) character error rate (CER) computed from Whisper-large-v3 transcriptions and (ii) human preference aggregated into ELO scores. This ELO ranking is derived from 596 total votes from 27 participants, who each provided pairwise blind judgments. Aggregate results are summarized in Table 1, with ELO rankings visualized in Figure 1.
@@ 5 Conclusion
Without a verifiable, automatic reward for prosody, GRPO trained on transcription-centric signals (CER/Whisper-NLL) predictably optimizes what is measured—intelligibility—while collapsing what is not—prosodic variation—into near-monotone speech. Extending the reward with speaker-similarity injects noisy, non-prosodic supervision that destabilizes optimization (e.g., EOS failures) and inflates CER, indicating that the core limitation lies in the reward design, not the optimizer.
We close this reward gap with small-batch iterative DPO, replacing unverifiable proxies with directly verifiable human preferences. With only 200 preference pairs per round, DPO consistently restores prosodic diversity favored by listeners (highest ELO) while keeping CER competitive, serving as a data-efficient complement to GRPO. We also release KoCC-TTS for robustness and conversational prosody evaluation. Our takeaway is simple: when prosody cannot be reliably rewarded automatically, human-in-the-loop preference optimization is the practical path to natural and robust TTS.

### v2

@@ 1 Introduction
To close this reward gap, we adopt iterative Direct Preference Optimization (DPO) with small human-in-the-loop batches. Across rounds, DPO supplies a directly verifiable signal for prosodic naturalness while regularizing to the current model, yielding both improved human preference and competitive CER (Fig. 1). Our contributions are as follows:
@@
  • We release KoCC-TTS, an evaluation set of frequent Korean call-center phrases for TTS prosody assessment. [^1] https://huggingface.co/datasets/channelcorp/KoCC-TTS-testset
@@ 2 Related Work
GRPO for TTS and reward design Recent TTS studies that adopt group-relative policy optimization (GRPO) primarily target intelligibility and identity preservation by rewarding ASR-derived errors and speaker similarity, sometimes adding non-intrusive quality predictors. For instance, F5R-TTS couples WER with speaker-similarity (SIM) under GRPO [15], DMOSpeech2 optimizes a duration policy with SIM+WER via GRPO [8], and the TTS-1 technical report describes a composite GRPO reward that blends WER/SIM/DNSMOS for RL alignment [1]. While these report lower error rates and stronger speaker faithfulness, they largely omit explicit prosody-sensitive rewards (e.g., pitch movement, phrasing, boundary control). In practice, we also observed that CER/NLL oriented GRPO can collapse prosodic variation into near-monotone renderings, consistent with reports of punctuation and phrasing related failures in state-of-the-art systems [14].
Preference based objectives for prosody Direct Preference Optimization (DPO) offers a complementary route by optimizing pairwise human (or proxy) preferences without training a separate reward model [11]. In TTS, Emo-DPO applies DPO to better capture subtle emotional/prosodic nuances with an LLM-based TTS backbone, improving both prosody similarity and perceived naturalness [6]. Beyond single shot DPO, iterative preference optimization for speech generation has also been explored. SpeechAlign constructs codec-token preference pairs and refines a speech LM in multiple rounds, demonstrating iterative self-improvement [17]. Concurrently, differentiable or multi-dimensional preference objectives have been proposed to move past coarse ASR metrics. DiffRO directly optimizes differentiable rewards over codec tokens [5], and MPO considers multi-criteria screening of preference pairs in speech synthesis [18]. These studies collectively suggest that preference based post training is a promising way to recover communicative prosody without sacrificing robustness to transcription errors.
@@ 4.2 Setup
We evaluate 12 systems on the KoCC-TTS dataset, including 3 production-grade external services, 3 open-source models, and 6 internal variants. We intentionally exclude open-source TTS baselines from the main comparison. A preliminary screening indicated that most off-the-shelf OSS voices exhibited inadequate prosodic fluency in Korean, and their inclusion would likely result in floor effects rather than provide a meaningful basis for comparison. For external systems, we adopt vendors’ strongest Korean voices and default synthesis settings unless otherwise noted: ElevenLabs Multilingual v2 (”Anna”) [^4] https://elevenlabs.io/blog/eleven-multilingual-v2, Supertone (”sona_speech_1”) [^5] https://docs.supertoneapi.com/ko/user-guide/quickstart, and GPT-4o-mini-tts (”sage”). To ensure fairness, all systems synthesize from the same prompts with identical text normalization rules. Speaking rate and punctuation handling are held fixed, and outputs are evaluated at vendors’ native sampling configurations.
We report (i) character error rate (CER) computed from Whisper-large-v3 transcriptions and (ii) human preference aggregated into ELO scores. For human evaluation, we adopt blind A/B pairwise comparison following Chatbot Arena-style evaluation [3]. We collected 596 votes from 27 participants whose ages ranged from 20 to 60. In each trial, raters listened to two anonymized audio samples and selected win, loss, or tie based on which utterance sounded more natural in terms of pitch and prosodic flow. Votes were aggregated via ELO-style ranking. Aggregate results are summarized in Table 1, with ELO rankings visualized in Figure 1.
@@ 4.5 Iterative DPO: Small Preference Sets Recover Prosody & CER
Why does Round 2 peak? We hypothesize that early rounds benefit from larger reward gaps between chosen and rejected samples, providing more informative gradients for preference learning. As iterations proceed, the policy-reference gap narrows and the marginal informativeness of new preference pairs diminishes, leading to saturation. This pattern is consistent with diminishing returns observed in iterative preference optimization [10].
@@ 5 Conclusion
Without a verifiable automatic reward for prosody, GRPO trained on transcription-centric signals (CER/Whisper-NLL) predictably optimizes what is measured—intelligibility—while collapsing what is not—prosodic variation—into near-monotone speech. Extending the reward with speaker-similarity injects noisy, non-prosodic supervision that destabilizes optimization (e.g., EOS failures) and inflates CER, indicating that the core limitation lies in the reward design, not the optimizer.
We close this reward gap with iterative DPO, replacing unverifiable proxies with directly verifiable human preferences. With only 200 preference pairs per round, DPO consistently restores prosodic diversity favored by listeners (highest ELO) while keeping CER competitive, serving as a data-efficient complement to GRPO. We also release KoCC-TTS for robustness and conversational prosody evaluation. Our takeaway is simple: when prosody cannot be reliably rewarded automatically, human-in-the-loop preference optimization is the practical path to natural and robust TTS.